You are listening to the radio when one of your favorite songs comes on, so you turn up the volume. If you increase the sound intensity by 29.0 dB, by what ratio did the intensity of the sound, in W/m2, increase? The intensity of sound, in W/m2, increased by a factor of
Answer #1

Since , sound level(L) = 10*log(I/Io) dB

So,

L2 - L1 = 10*log(I2/Io) - 10*log(I1/Io)

where, Io = 10^-12 W/m^2

I1 & I2 = sound intensity

given , L2 - L1 = 29.0 dB

So, 29.0 = 10*log(I2/Io) - 10*log(I1/Io)

29.0 = 10*log[(I2/Io)/(I1/Io)]

2.9 = log(I2/I1)

I2/I1 = 10^2.9

I2/I1 = 794

Therefore intensity of sound is increased by 794.

Please upvote.